1/4=3-2x+-1/x+2
Can you please solve the solution to this alegbra equation.

Answer by ad_alta(240)

1) 1/4=3-2x+-1/x+2 [subtract 5 from both sides]
2) -19/4=-2x-1/x [common denominator]
3) -19/4=(-2x^2-1)/x [quadratic]
4) x=(19 {+/-} sqrt(233))/16